Suppose that the equation p(x) = -5x^2 + 600x +100 (where x represents the number of items sold), describes the profit function for a certain business. How many items should be sold to maximize the profit?
Answer by ewatrrr(24785)

 
Hi,
p(x) = -5x^2 + 600x +100 =
p(x) = -5(x-60)^2 + 18100 |Completing the Square, Vmax = (60,18100)
at 60 units sold, profit is at a maximum